OpenStudy (anonymous):

Britney throws an object into the air with an initial vertical velocity of 27 feet per second from a platform that is 10 feet above the ground. About how long will it take the object to hit the ground? A. 1 second B. 2 seconds C. 3 seconds D. 4 seconds

OpenStudy (anonymous):

@g17 @mathstudent55 @~Jane_Doe~

OpenStudy (anonymous):

@whpalmer4

OpenStudy (anonymous):

did you come up with an equation?

OpenStudy (whpalmer4):

the formula for the height is \[h(t) = -\frac{1}{2}gt^2 + v_0t + h_0\]where \[g = 32\frac{\text{ft}}{\text{s}^2}\]\[h_0 =10\text{ ft}\]\[v_0=27\frac{\text{ ft}}{\text{ s}}\]

OpenStudy (whpalmer4):

If you solve that for the value of \(t\) that makes \(h(t) = 0\) and makes sense in this context, you'll have your answer.

OpenStudy (anonymous):

ok thanks :)

OpenStudy (whpalmer4):

The height of the object is a parabola if plotted with time as the x-axis and height as the y-axis. The vertex of the parabola is the maximum height it reaches, and the positive solution is where the parabola crosses the x-axis with a positive value of x.